OpenStudy (anonymous):

Which of the following options is an equivalent function to f(x) = 4(3)2x? f(x) = 36x f(x) = 4(9)x f(x) = 144x f(x) = 4(6x)

TheSmartOne (thesmartone):

Is your equation actually \(\sf\Large f(x) = 4(3)^{2x}\) Or is it \(\sf\Large f(x) = 4(3)(2x)\)

OpenStudy (anonymous):

its the first one

OpenStudy (anonymous):

@TheSmartOne please help me

TheSmartOne (thesmartone):

ok, so we can say that x + x = 2x right?

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

and do you know this property \(\sf\Large a^b \times a^b = a^{b+b}\)

OpenStudy (anonymous):

kind of

TheSmartOne (thesmartone):

do you know the property: \(\sf\Large (a^b)^c = a^{bc}\)

OpenStudy (anonymous):



TheSmartOne (thesmartone):

ok, good, we're going to use it :)

TheSmartOne (thesmartone):

\(\sf\Large f(x) = (4)(3^{2x}) = (4)(3^2)^x\) What is 3^2 = 3 x 3 = ?

TheSmartOne (thesmartone):

@Mal4902